OpenStudy (anonymous):

I really need help with Algebra 2 Quadratic Relations and Conic Functions! Identify the conic section, any lines of symmetry, and the domain and range. x=2y^2+5 I need help understand how to do this! I will fan anyone who can help me! 😰

OpenStudy (anonymous):

umm... I'm not sure. I'll fan you to make up for my failure to help you. :P

OpenStudy (anonymous):

@tom982

OpenStudy (anonymous):

Firstly, do you know how to identify a conic section?

OpenStudy (anonymous):

A circle is when x^2 and y^2 have the same coefficients, an ellipse is when it's the same sign but different coefficients, and a hyperbola is when it's different signs.

OpenStudy (anonymous):

@tom982

OpenStudy (anonymous):

@tom982 please help me :(

OpenStudy (anonymous):

@CGGURUMANJUNATH

OpenStudy (anonymous):

Is that meant to be x^2 in the original equation then? You're talking about x^2 when it's not relevant here.

OpenStudy (anonymous):

It says the equation is "x=2y^2+5" and I need to find the conic section, lines of symmetry, and the domain and range @tom982

OpenStudy (anonymous):

Okay. Then you're barking up the wrong tree with x^2. Read this: http://www.ck12.org/book/CK-12-Algebra-II-with-Trigonometry-Concepts/section/10.11/ Can you classify it now?

OpenStudy (anonymous):

Hyperbola?

OpenStudy (anonymous):

A conic is of the form \(Ax^2+Bxy+Cy^2+Dx+Ey+F=0\). Rearranging yours into this form we get: \(2y^2-x+5=0\). This gives us our constants \(A=0\), \(B=0\), \(C=2\)\, \(D=-1\), \(E=0\) and \(F=5\). The discriminant is \(B^2-4AC\) and in this case it is \(0^2-4\times0\times2=0\). As the discriminant is zero, what does this say about our conic? Refer to the page I linked you to.

OpenStudy (anonymous):

Ah so a parabola

OpenStudy (anonymous):

Exactly, well done. Parabolas are symmetric about the x-axis whenever (x,y)=(x,-y) which we can easily check: \(x=2(-y)^2+5=2y^2+5\) as required. Hence the x-axis (the line y=0) is the only line of symmetry for this parabola. Can you find the domain and range?

OpenStudy (anonymous):

Domain: -5<=x<=5 Range: -5<=y<=5?

OpenStudy (anonymous):

Why have you restricted the domain? What's wrong with using -6? \(x=2(-6)^2+5=77\)

OpenStudy (anonymous):

Wait how did you get that???

OpenStudy (anonymous):

Ah sorry I misread what you said, but you're still wrong. Your function is written abnormally: \(f(y)=x\) so our domain is for y and the range is for x. There are no numbers we can't put into this equation so the domain is \(y \in \mathbb{R} \). What numbers can we get out of it?